OpenStudy (anonymous):

4-5/8 X 4 how does this come to 3/2

OpenStudy (anonymous):

(32/8)-[(5/8)*4] (32/8) - (5/2) (32/8) - (20/8) which equals 12/8 reduce down to 3/2

OpenStudy (anonymous):

so do i subtract first

OpenStudy (hoblos):

no you really have to do multiplication first so 4 - 5/8 x 4 = 4 - 5/2 = 8/2 - 5/2 =3/2 :D